OpenStudy (anonymous):

Write: 3log(x+4) - 2log(x-7) + 5log(x-2) - log(x^2) as a single logarithm.

OpenStudy (anonymous):

\[alogx= logx^a\] \[\log a + \log b= \log ab \]\[\log a - \log b= loga/b\]

OpenStudy (anonymous):